# Structure-Based Molecule Optimization via Gradient-Guided Bayesian Update

### v2

@@ Abstract
Structure-based molecule optimization (SBMO) aims to optimize molecules with both continuous coordinates and discrete types against protein targets. A promising direction is to exert gradient guidance on generative models given its remarkable success in images, but it is challenging to guide discrete data and risks inconsistencies between modalities. To this end, we leverage a continuous and differentiable space derived through Bayesian inference, presenting Molecule Joint Optimization (MolJO), the first gradient-based SBMO framework that facilitates joint guidance signals across different modalities while preserving SE(3)-equivariance. We introduce a novel backward correction strategy that optimizes within a sliding window of the past histories, allowing for a seamless trade-off between explore-and-exploit during optimization. Our proposed MolJO achieves state-of-the-art performance on CrossDocked2020 benchmark (Success Rate 51.3%, Vina Dock -9.05 and SA 0.78), more than 4×4$\times$4 × improvement in Success Rate compared to the gradient-based counterpart, and 2×2$\times$2 × “Me-Better” Ratio as much as 3D baselines. Furthermore, we extend MolJO to a wide range of optimization settings, including multi-objective optimization and challenging tasks in drug design such as R-group optimization and scaffold hopping, further underscoring its versatility and potential.
@@ 1 Introduction
Structure-based drug design (SBDD) plays a critical role in drug discovery by identifying three-dimensional (3D) molecules that are favorable against protein targets (Isert et al., 2023). While recent SBDD focuses on the initial identification of potential drug candidates, these compounds must undergo a series of further modifications for optimized properties, a process that is both complex and time-consuming (Hughes et al., 2011). Therefore, structure-based molecule optimization (SBMO) has garnered increasing interest in real-world drug design (Zhou et al., 2024), emphasizing the practical need of optimizing 3D molecules to meet specific therapeutic criteria.
@@
Concretely, SBMO can be viewed as a more advanced task within the broader scope of general SBDD, requiring precise control over molecular properties while navigating the chemical space. Specifically, SBMO addresses two key aspects: (1) SBMO prioritizes targeted molecular property enhancement according to expert-specified objectives, whereas generative models for SBDD primarily focus on maximizing data likelihood without special emphasis on property improvement (Luo et al., 2021; Peng et al., 2022). Therefore these models can only produce outputs similar to their training data, limiting the ability to improve molecular properties. (2) SBMO is capable of optimizing existing compounds with 3D structural awareness, addressing a critical gap left by previous molecule optimization methods with 1D SMILES or 2D graph representations (Bilodeau et al., 2022; Fu et al., 2022), and allowing for a more nuanced control. The focus on structure makes SBMO particularly suited for key design tasks, such as R-group optimization and scaffold hopping.
A pioneering work for SBMO is DecompOpt (Zhou et al., 2024), which designs a special 3D generative model conditioned on fragments, and employs a gradient-free sampling method within such decomposed fragment space through iterative oracle calls. One of the drawbacks is that it relies on oracle functions that can be computationally expensive especially in large-scale SBMO tasks, where running multiple rounds of costly simulations is impractical or even infeasible. Moreover, DecompOpt’s contribution lies in its novel generative paradigm based on fragment conditions, rather than exploring optimization within existing models, making it less adaptable to other frameworks.
@@
One generalized solution is gradient guidance, addressing these issues by eliminating the need for expensive oracle simulations while being flexible enough to be incorporated into strong-performing generative models in a plug-and-play fashion, as demonstrated in a wide range of challenging real-world applications including image synthesis (Dhariwal & Nichol, 2021; Epstein et al., 2023).
However, current gradient-based methods have not fully realized their potential in SBMO, for they have historically suffered from the continuous-discrete challenges: (1) it is non-trivial to guide discrete variable within probabilistic generative process. More specifically, standard gradient guidance is designed for continuous variables that follow Gaussian distributions, making them not directly applicable to molecular data that involve discrete atom types. Methods attempting to adapt gradient guidance to discrete data often resort to approximating these variables as continuous, either by adding Gaussian noise (Bao et al., 2022) or by assuming that classifiers follow a Gaussian distribution (Vignac et al., 2023). Unfortunately, these approximations can lead to suboptimal results, as they do not accurately reflect the discrete nature (Kong et al., 2023). (2) Gradient guidance might introduce inconsistencies between modalities. For instance, TAGMol (Dorna et al., 2024) formalizes guidance exclusively over continuous coordinates, resulting in a disconnect between the discrete and continuous modalities. This may explain why TAGMol struggles to optimize overall molecular properties as shown in Figure 1, despite its improvement in Vina affinities. By solely guiding the continuous coordinates, TAGMol enhances spatial protein-ligand interactions but fails to optimize e.g. synthesizeability, which depends more on molecular topology, especially discrete atom types.
In this paper, we address the multi-modality challenge for gradient guidance by leveraging a continuous and differential space, representing an aggregation of noisy samples from the data space derived through Bayesian inference (Graves et al., 2023). We design MolJO, a general framework for Joint Optimization of 3D molecules and the first proof-of-concept for gradient-based optimization of continuous-discrete variables. We introduce a novel sampling strategy called backward correction, enhancing the alignment of gradients over different steps. By maintaining a sliding window of past history for optimization, the backward correction strategy enforces explicit dependency on the past, effectively alleviating the issue of inconsistencies. Moreover, it balances the exploration of molecular space with the exploitation of better-aligned guidance signals, offering a flexible trade-off.
@@
Figure 1: Overview. A. Structure-based molecule optimization, including (1) guiding molecule design by expert-specified objectives, (2) optimizing existing compounds in the structure space. B. Study on the ratio of “me-better” molecules, where all other baselines fall short in the overall improvement. C. Overall illustration of MolJO, utilizing joint gradient signals over continuous-discrete data, where the distributions of $\bm{\theta}$ (continuous $\bm{\mu}$ and discrete $\mathbf{z}$) are taken from true guided trajectories. D. Graphical model of our proposed backward correction strategy, keeping a sliding window of size $k$.
Our main contributions are summarized as follows:
@@
  • We propose MolJO, the first joint gradient-based method for SBMO that establishes the guidance over multi-modality molecular data, offering better controllability and paving the way for gradient guidance over continuous-discrete variables.
@@
  • MolJO achieves the best Vina Dock of -9.05, SA of 0.78 and Success Rate of 51.3%, and “Me-Better” Ratio of improved molecules that is $\bf 2\times$ as much as other 3D baselines. We generalize MolJO to various needs including R-group optimization and scaffold hopping, highlighting its versatility and potential in practical drug design.
@@ 2 Related Work
Pocket-aware generative models aim to learn a conditional distribution over the protein-ligand complex data. Initial approaches adopt 1D SMILES or 2D graph representation (Bjerrum & Threlfall, 2017; Gómez-Bombarelli et al., 2018), and recent research has shifted its focus towards 3D molecule generation in order to better capture the interatomic interactions. Early atom-based autoregressive models (Luo et al., 2021; Peng et al., 2022; Liu et al., 2022) enforce an atom ordering to generate molecules atom-by-atom. Fragment-based methods (Powers et al., 2022; Zhang et al., 2023; Lin et al., 2023) alleviate the issue of ordering by decomposing molecules into motifs instead of atom-level generation, but they risk more severe error accumulation and thus generally require post-processing or multi-stage treatment. Non-autoregressive methods based on diffusions (Schneuing et al., 2022; Guan et al., 2022; 2023) and BFNs (Qu et al., 2024) target full-atom generation for enhanced performance and efficiency. However, the needs of optimizing certain properties and modifying existing compounds are not adequately addressed in the scope of previous methods, limiting their usefulness in drug design.
@@
Inspired by classifier guidance for diffusions (Dhariwal & Nichol, 2021), pioneering approaches are committed to adapting the guidance method to handle the complicated molecular geometries in the setting of pocket-unaware generation. EEGSDE (Bao et al., 2022) derives an equivariant framework for continuous diffusion, and MUDM (Han et al., 2023) further explores time-independent property functions for guidance. As they enforce a continuous diffusion process for discrete variables, these methods are not applicable in advanced molecular modeling (Guan et al., 2022; 2023) where discrete data are processed by a discrete diffusion, for it is unnatural to apply progressive Gaussian noise that drives Categorical data away from the simplex. DiGress (Vignac et al., 2023) proposes classifier guidance for discrete diffusion of molecular graphs, yet it additionally assumes that the probability of classifier follows a Gaussian, which is ungrounded and often a problematic approximation. Based on the continuous-discrete diffusion for SBDD, TAGMol (Dorna et al., 2024) retains the guidance only for continuous coordinates, because there lacks a proper way to propagate the gradient over discrete types. The discrete part is affected only implicitly and belatedly in the generative process, and such imbalanced guidance would probably result in suboptimal performance for lack of joint optimization.
@@ 3 Preliminary
A protein binding site $\mathbf{p}=(\mathbf{x}_{P},\mathbf{v}_{P})$ is represented as a point cloud of $N_{P}$ atoms where atom coordinates $\mathbf{x}_{P}=\{\mathbf{x}_{P}^{1},\dots,\mathbf{x}_{P}^{N_{P}}\}\in\mathbb{R}^{N_{P}\times 3}$ and $K_{P}$-dimensional atom features $\mathbf{v}_{P}=\{\mathbf{v}_{P}^{1},\dots,\mathbf{v}_{P}^{N_{P}}\}\in\mathbb{R}^{N_{P}K_{P}}$. Similarly, a ligand molecule $\mathbf{m}=(\mathbf{x}_{M},\mathbf{v}_{M})$ contains $N_{M}$ atoms, where $\mathbf{x}_{M}^{(i)}\in\mathbb{R}^{3}$ is the atomic coordinate and $\mathbf{v}_{M}^{(i)}\in\mathbb{R}^{K_{M}}$ the atom type. For brevity, the subscript for molecules $\cdot_{M}$ and the pocket condition $\mathbf{p}$ are omitted unless necessary.
@@
A key motivation for BFN is that the transmission ought to be continuous and smooth, therefore it does not directly operate on the noisy latent $\mathbf{y}$ as diffusions, but on the structured Bayesian posterior $\bm{\theta}$ given noisy latents instead. The receiver holds a prior belief $\bm{\theta}_{0}$, and updates the belief by aggregating noisy observations $\mathbf{y}$, yielding the Bayesian update distribution:
$p_{U}(\bm{\theta}_{i}~{}|~{}\bm{\theta}_{i-1},\mathbf{x};\alpha_{i})=\underset{p_{S}(\mathbf{y}_{i}|\mathbf{x};\alpha_{i})}{\mathbb{E}}\delta\Big{(}\bm{\theta}_{i}-h(\bm{\theta}_{i-1},\mathbf{y}_{i},\alpha_{i})\Big{)}$ (1)
where $\delta(\cdot)$ is Dirac distribution, and Bayesian update function $h$ is derived through Bayesian inference. Note that $\bm{\theta}_{i}$ can be viewed as the result of mapping given the sender distributions of $\mathbf{y}_{0}$ to $\mathbf{y}_{i}$:
$\bm{\theta}_{i}=f(\mathbf{y}_{0},\mathbf{y}_{1},\dots,\mathbf{y}_{i})$ (2)
@@
$L^{n}(\mathbf{x})=\underset{\prod_{i=1}^{n}p_{U}(\bm{\theta}_{i}|\bm{\theta}_{i-1},\mathbf{x};\alpha_{i})}{\mathbb{E}}\sum_{i=1}^{n}D_{\text{KL}}(p_{S}(\mathbf{y}_{i}~{}|~{}\mathbf{x};\alpha_{i})~{}\|~{}p_{R}(\mathbf{y}_{i}~{}|~{}\bm{\theta}_{i-1},t_{i},\alpha_{i})).$ (4)
@@ 4 Method
In this section, we introduce MolJO that guides the distribution over $\bm{\theta}$, utilizing aggregated information from previous latents. Though different from guided diffusions that operate on noisy latent $\mathbf{y}$, this guidance aligns with our generative process conditioned on $\bm{\theta}$. By focusing on $\bm{\theta}$, we can effectively steer the clean samples towards desirable direction, ensuring a smooth gradient flow.
@@
As illustrated in Figure 1, we introduce MolJO as follows: in Sec. 4.1, we propose the concept of gradient guidance over the multi-modality molecule space, derive the form of guided transition kernel $\pi(\bm{\theta}_{i}|\bm{\theta}_{i-1})$ via first-order Taylor expansion, and explain the underlying manipulations of distributions the guidance corresponds to. In Sec. 4.2, we present a generalized advanced sampling strategy termed backward correction for $p_{\phi}$, which allows for a flexible trade-off between explore-and-exploit by maintaining a sliding window of past histories. We empirically demonstrate our strategy helps optimize consistency across steps, ultimately improving the overall performance.
@@ 4.1 Equivariant Guidance for multi-modality Molecular Data
In this section, we derive the detailed guidance over $\bm{\theta}$ for molecule $\mathbf{m}=(\mathbf{x},\mathbf{v})$ with $N$ atoms, where $\mathbf{x}\in\mathbb{R}^{N\times 3}$ represent continuous atom coordinates and $\mathbf{v}\in\{1,\dots,K\}^{N}$ for $K$ discrete atom types, and thus $\bm{\theta}:=[\bm{\mu},\mathbf{z}]$ for the continuous and discrete modality, respectively.
@@
Assuming $\bm{\mu}_{i}\sim\mathcal{N}(\bm{\mu}_{\phi},\sigma)$ and $\mathbf{y}_{i}\sim\mathcal{N}(\mathbf{y}_{\phi},\sigma^{\prime})$ in the original generative process of BFN, we can approximate the guided transition kernel $\pi(\bm{\theta}_{i}|\bm{\theta}_{i-1})$ by sampling
$\bm{\mu}_{i}^{*}$ $\sim\mathcal{N}(\bm{\mu}_{\phi}+\sigma\mathbf{g}_{\bm{\mu}},\sigma)$ (5)
$\mathbf{y}_{i}^{*}$ $\sim\mathcal{N}(\mathbf{y}_{\phi}+\sigma^{\prime}\mathbf{g}_{\mathbf{y}},\sigma^{\prime})$ (6)
where gradient $\mathbf{g}_{\bm{\mu}}=-\nabla_{\bm{\mu}}E(\bm{\theta},t_{i})|_{\bm{\theta}=\bm{\theta}_{i-1}}$, $\mathbf{g}_{\mathbf{y}}=-\nabla_{\mathbf{y}}E(\bm{\theta},t_{i})|_{\bm{\theta}=\bm{\theta}_{i-1}}$, recall $\bm{\theta}:=[\bm{\mu},\mathbf{z}=f(\mathbf{y})]$.
@@
Guiding $\bm{\mu}$ for Continuous $\mathbf{x}$.
For continuous coordinates $\mathbf{x}\in\mathbb{R}^{N\times 3}$, it is natural to adopt a Gaussian sender distribution $\mathbf{y}^{x}\sim\mathcal{N}(\mathbf{x},\alpha^{-1}\mathbb{I})$. With a prior belief $\bm{\mu}_{0}=\bm{0}$, we have the Bayesian update function for posterior $\bm{\mu}_{i}$ given noisy $\mathbf{y}^{x}$ as in Graves et al. (2023):
$h(\bm{\mu}_{i-1},\mathbf{y}^{x},\alpha_{i})=\frac{\bm{\mu}_{i-1}\rho_{i-1}+\mathbf{y}^{x}\alpha_{i}}{\rho_{i}}$ (7)
with $\alpha_{i}=\beta_{i}-\beta_{i-1},\rho_{i}=1+\beta_{i}$ given the schedule $\beta_{i}=\sigma_{1}^{-2i/n}-1$ for a positive $\sigma_{1}$ and $n$ steps.

### v3

# Empower Structure-Based Molecule Optimization with Gradient Guidance

@@ Abstract
Structure-based molecule optimization (SBMO) aims to optimize molecules with both continuous coordinates and discrete types against protein targets. A promising direction is to exert gradient guidance on generative models given its remarkable success in images, but it is challenging to guide discrete data and risks inconsistencies between modalities. To this end, we leverage a continuous and differentiable space derived through Bayesian inference, presenting Molecule Joint Optimization (MolJO), the gradient-based SBMO framework that facilitates joint guidance signals across different modalities while preserving SE(3)-equivariance. We introduce a novel backward correction strategy that optimizes within a sliding window of the past histories, allowing for a seamless trade-off between explore-and-exploit during optimization. MolJO achieves state-of-the-art performance on CrossDocked2020 benchmark (Success Rate 51.3%, Vina Dock -9.05 and SA 0.78), more than 4×4$\times$4 × improvement in Success Rate compared to the gradient-based counterpart, and 2×2$\times$2 × “Me-Better” Ratio as much as 3D baselines. Furthermore, we extend MolJO to a wide range of optimization settings, including multi-objective optimization and challenging tasks in drug design such as R-group optimization and scaffold hopping, further underscoring its versatility.
Bayesian Flow Network, Structure-Based Molecule Optimization
[…]
@@ 1 Introduction
Structure-based drug design (SBDD) plays a critical role in drug discovery by identifying three-dimensional (3D) molecules that are favorable to protein targets (Isert et al., 2023). While recent SBDD focuses on the initial identification of potential drug candidates, these compounds must undergo a series of further modifications for optimized properties, a process that is both complex and time-consuming (Hughes et al., 2011). Therefore, structure-based molecule optimization (SBMO) has garnered increasing interest in real-world drug design (Zhou et al., 2024a), emphasizing the practical need to optimize 3D molecules to meet specific therapeutic criteria.
@@
Concretely, SBMO can be viewed as a more advanced task within the broader scope of general SBDD, requiring precise control over molecular properties while navigating the chemical space. Specifically, SBMO addresses two key aspects: (1) SBMO prioritizes targeted molecular property enhancement according to expert-specified objectives, while generative models for SBDD focus primarily on maximizing the likelihood of data without special emphasis on property improvement (Luo et al., 2021; Peng et al., 2022). Therefore, these models can only produce outputs similar to their training data, limiting the ability to improve molecular properties. (2) SBMO is capable of optimizing existing compounds with 3D structural awareness, addressing a critical gap left by previous molecule optimization methods with 1D SMILES or 2D graph representations (Bilodeau et al., 2022; Fu et al., 2022), and allowing for a more nuanced control. The focus on structure makes SBMO particularly suited for key design tasks, such as R-group optimization and scaffold hopping.
@@
Recent works have explored SBMO through evolutionary-based resampling and gradient-based methods. For instance, DiffSBDD (Schneuing et al., 2022) uses a gradient-free evolutionary sampling method, and DecompOpt (Zhou et al., 2024a) further introduces a fragment-conditioned 3D generative model for resampling. These approaches rely on iterative, computationally expensive oracle calls to select top-of-N candidates. One generalized and orthogonal solution is gradient guidance, eliminating the need for oracle simulations while being flexible enough to be incorporated into strong-performing generative models in a plug-and-play fashion, as demonstrated in a wide range of challenging real-world applications including image synthesis (Dhariwal & Nichol, 2021; Epstein et al., 2023).
@@
Figure 1: Overview. A. Structure-based molecule optimization, including (1) guiding molecule design by expert-specified objectives, (2) optimizing existing compounds in the structure space. B. Study on the ratio of “me-better” molecules (with improved properties), where all other baselines fall short in the overall improvement. C. Overall illustration of MolJO, utilizing joint gradient signals over continuous-discrete data, where the distributions of $\bm{\theta}$ (for continuous $\mathbf{x}$ and discrete $\mathbf{v}$) are taken from true guided trajectories. D. Graphical model of our proposed backward correction strategy, keeping a sliding window of size $k$.
However, current gradient-based methods have not fully realized their potential in SBMO, for they have historically suffered from the continuous-discrete challenges: (1) it is non-trivial to guide discrete variable within probabilistic generative process. More specifically, standard gradient guidance is designed for continuous variables that follow Gaussian distributions, making them not directly applicable to molecular data that involve discrete atom types. Methods attempting to adapt gradient guidance to discrete data often resort to approximating these variables as continuous, either by adding Gaussian noise (Bao et al., 2022) or by assuming that classifiers follow a Gaussian distribution (Vignac et al., 2023). Unfortunately, these approximations can lead to suboptimal results, as they do not accurately reflect the discrete nature (Kong et al., 2023). (2) Gradient guidance might introduce inconsistencies between modalities. For instance, TAGMol (Dorna et al., 2024) formalizes guidance exclusively over continuous coordinates, resulting in a disconnect between the discrete and continuous modalities. This may explain why TAGMol struggles to optimize overall molecular properties as shown in Fig. 1, despite its improvement in Vina affinities. By solely guiding the continuous coordinates, TAGMol enhances spatial protein-ligand interactions but fails to optimize e.g. synthesizeability, which depends more on molecular topology, especially discrete atom types.
In this paper, we address the multi-modality challenge for gradient guidance by leveraging a continuous and differential space, representing an aggregation of noisy samples from the data space derived through Bayesian inference (Graves et al., 2023). We design MolJO (Molecule Joint Optimization), a principled, ent-to-end differentiable framework that enables gradient-based optimization of continuous and discrete variables. We introduce a novel sampling strategy called backward correction, enhancing the alignment of gradients over different steps. By maintaining a sliding window of past history for optimization, the backward correction strategy enforces explicit dependency on the past, effectively alleviating the issue of inconsistencies. Moreover, it balances the exploration of molecular space with the exploitation of better-aligned guidance signals, offering a flexible trade-off.
Our main contributions are summarized as follows.
@@
  • We propose MolJO, the joint gradient-based method for SBMO that establishes the guidance over molecules, offering better controllability and effectively integrating gradient guidance for continuous-discrete variables within a unified framework.
@@
  • MolJO achieves the best Vina Dock of -9.05, SA of 0.78 and Success Rate of 51.3%, and “Me-Better” Ratio of improved molecules that is $\bf 2\times$ as much as other 3D baselines. We generalize MolJO to various needs including R-group optimization and scaffold hopping, highlighting its versatility.
@@ 2 Related Work
Pocket-aware generative models aim to learn a conditional distribution over the protein-ligand complex data. Initial approaches adopt 1D SMILES or 2D graph representation (Bjerrum & Threlfall, 2017; Gómez-Bombarelli et al., 2018), and recent research has shifted its focus towards 3D molecule generation in order to better capture interatomic interactions. Early atom-based autoregressive models (Luo et al., 2021; Peng et al., 2022; Liu et al., 2022) enforce an atom ordering to generate molecules atom-by-atom. Fragment-based methods (Powers et al., 2022; Zhang et al., 2023; Lin et al., 2023) alleviate the issue of ordering by decomposing molecules into motifs instead of atom-level generation, but they risk more severe error accumulation and thus generally require post-processing or multi-stage treatment. Non-autoregressive methods based on diffusions (Schneuing et al., 2022; Guan et al., 2022, 2023) and BFNs (Qu et al., 2024) target full-atom generation for enhanced performance and efficiency. However, the needs of optimizing certain properties and modifying existing compounds are not adequately addressed in the scope of previous methods, limiting their usefulness in drug design.
@@
Inspired by classifier guidance for diffusions (Dhariwal & Nichol, 2021), pioneering approaches are committed to adapting the guidance method to handle the complicated molecular geometries in the setting of pocket-unaware generation. EEGSDE (Bao et al., 2022) derives an equivariant framework for continuous diffusion, and MUDM (Han et al., 2023) further explores time-independent property functions for guidance. As they enforce a continuous diffusion process for discrete variables, these methods are not applicable in advanced molecular modeling (Guan et al., 2022, 2023) where discrete data are processed by a discrete diffusion, for it is unnatural to apply progressive Gaussian noise that drives Categorical data away from the simplex. DiGress (Vignac et al., 2023) proposes classifier guidance for discrete diffusion of molecular graphs, yet it additionally assumes that the probability of classifier follows a Gaussian, which is ungrounded and often a problematic approximation. Based on the continuous-discrete diffusion for SBDD, TAGMol (Dorna et al., 2024) retains the guidance only for continuous coordinates, because there lacks a proper way to propagate the gradient over discrete types. The discrete part is affected only implicitly and belatedly in the generative process, and such imbalanced guidance would probably result in suboptimal performance for lack of joint optimization.
@@ 3 Preliminary
A protein binding site $\mathbf{p}=(\mathbf{x}_{P},\mathbf{v}_{P})$ is represented as a point cloud of $N_{P}$ atoms with coordinates $\mathbf{x}_{P}=\{\mathbf{x}_{P}^{1},\dots,\mathbf{x}_{P}^{N_{P}}\}\in\mathbb{R}^{N_{P}\times 3}$ and $K_{P}$-dimensional atom features $\mathbf{v}_{P}=\{\mathbf{v}_{P}^{1},\dots,\mathbf{v}_{P}^{N_{P}}\}\in\mathbb{R}^{N_{P}K_{P}}$. Similarly, a ligand molecule $\mathbf{m}=(\mathbf{x}_{M},\mathbf{v}_{M})$ contains $N_{M}$ atoms, where $\mathbf{x}_{M}^{(i)}\in\mathbb{R}^{3}$ is the atomic coordinate and $\mathbf{v}_{M}^{(i)}\in\mathbb{R}^{K_{M}}$ the atom type. For brevity, the subscript for molecules $\cdot_{M}$ and the pocket condition $\mathbf{p}$ are omitted unless necessary.
@@
A key motivation for BFN is that the transmission ought to be continuous and smooth, therefore it does not directly operate on the noisy latent $\mathbf{y}$ as diffusions, but on the structured Bayesian posterior $\bm{\theta}$ given noisy latents instead. The receiver holds a prior belief $\bm{\theta}_{0}$, and updates the belief upon observed $\mathbf{y}$, yielding the Bayesian update distribution:
$\footnotesize p_{U}(\bm{\theta}_{i}~{}|~{}\bm{\theta}_{i-1},\mathbf{x};\alpha_{i})=\underset{p_{S}(\mathbf{y}_{i}|\mathbf{x};\alpha_{i})}{\mathbb{E}}\delta\Big{(}\bm{\theta}_{i}-h(\bm{\theta}_{i-1},\mathbf{y}_{i},\alpha_{i})\Big{)}$ (1)
where $\delta(\cdot)$ is Dirac distribution, and Bayesian update function $h$ is derived through Bayesian inference.
@@
$L^{n}(\mathbf{x})$ $=\underset{\prod_{i=1}^{n}p_{U}(\bm{\theta}_{i}|\bm{\theta}_{i-1},\mathbf{x};\alpha_{i})}{\mathbb{E}}$
$\sum_{i=1}^{n}D_{\text{KL}}$ $(p_{S}(\mathbf{y}_{i}~{}|~{}\mathbf{x};\alpha_{i})~{}\|~{}p_{R}(\mathbf{y}_{i}~{}|~{}\bm{\theta}_{i-1},t_{i},\alpha_{i})).$ (3)
@@ 4 Method
In this section, we introduce MolJO that guides the distribution over $\bm{\theta}$, utilizing aggregated information from previous latents. Though different from guided diffusions that operate on noisy latent $\mathbf{y}$, this guidance aligns with our generative process informed by $\bm{\theta}$. By focusing on lower-variance $\bm{\theta}$, we can effectively steer the clean samples towards desirable direction, ensuring a smooth gradient flow.
@@
As illustrated in Fig. 1, we introduce MolJO as follows: in Sec. 4.1, we propose the concept of gradient guidance over the multi-modality molecule space, derive the form of guided transition kernel $\pi(\bm{\theta}_{i}|\bm{\theta}_{i-1})$ via first-order Taylor expansion, and explain the underlying manipulations of distributions the guidance corresponds to. In Sec. 4.2, we present a generalized advanced sampling strategy termed backward correction for $p_{\phi}$, which allows for a flexible trade-off between explore-and-exploit by maintaining a sliding window of past histories. We empirically demonstrate our strategy helps optimize consistency across steps, ultimately improving the overall performance.
@@ 4.1 Equivariant Guidance for multi-modality Molecular Data
In this section, we derive the detailed guidance over $\bm{\theta}$ for molecule $\mathbf{m}=(\mathbf{x},\mathbf{v})$ with $N$ atoms, where $\mathbf{x}\in\mathbb{R}^{N\times 3}$ represent continuous atom coordinates and $\mathbf{v}\in\{1,\dots,K\}^{N}$ for $K$ discrete atom types, and thus $\bm{\theta}:=[\bm{\theta}^{x},\bm{\theta}^{v}]$, latent $\mathbf{y}:=[\mathbf{y}^{x},\mathbf{y}^{v}]$ for the continuous and discrete modality.
@@
Suppose $\tilde{\bm{\theta}}^{x}_{i}\sim\mathcal{N}(\bm{\theta}^{x}_{\phi},\sigma^{x}\mathbb{I})$ and $\tilde{\mathbf{y}}_{i}\sim\mathcal{N}(\mathbf{y}_{\phi},\sigma^{v}\mathbb{I})$ by definition of BFN generative process, we can approximate the guided transition kernel $\pi(\bm{\theta}_{i}|\bm{\theta}_{i-1})$:
$\bm{\theta}^{x}_{i}$ $\sim\mathcal{N}(\bm{\theta}^{x}_{\phi}+\sigma^{x}\mathbf{g}_{\bm{\theta}^{x}},\sigma^{x}\mathbb{I})$ (4)
$\mathbf{y}^{v}_{i}$ $\sim\mathcal{N}(\mathbf{y}^{v}_{\phi}+\sigma^{v}\mathbf{g}_{\mathbf{y}^{v}},\sigma^{v}\mathbb{I})$ (5)
where gradient $\mathbf{g}_{\bm{\theta}^{x}}=-\nabla_{\bm{\theta}^{x}}E(\bm{\theta},t_{i})|_{\bm{\theta}=\bm{\theta}_{i-1}}$, $\mathbf{g}_{\bm{\theta}^{v}}=-\nabla_{\bm{\theta}^{v}}E(\bm{\theta},t_{i})|_{\bm{\theta}=\bm{\theta}_{i-1}}$, $\mathbf{g}_{\mathbf{y}^{v}}=\mathbf{g}_{\bm{\theta}^{v}}\frac{\partial\bm{\theta}^{v}}{\partial\mathbf{y}^{v}}$.
@@
Guiding $\bm{\theta}^{x}$ for Continuous $\mathbf{x}$.
For continuous coordinates $\mathbf{x}\in\mathbb{R}^{N\times 3}$, it is natural to adopt a Gaussian sender distribution $\mathbf{y}^{x}\sim\mathcal{N}(\mathbf{x},\alpha^{-1}\mathbb{I})$. With a prior belief $\bm{\theta}^{x}_{0}=\bm{0}$, we have the Bayesian update function for posterior $\bm{\theta}^{x}_{i}$ given noisy $\mathbf{y}^{x}$ as in Graves et al. (2023):
$h(\bm{\theta}^{x}_{i-1},\mathbf{y}^{x},\alpha_{i})=\frac{\bm{\theta}^{x}_{i-1}\rho_{i-1}+\mathbf{y}^{x}\alpha_{i}}{\rho_{i}}$ (6)
with $\alpha_{i}=\beta^{x}_{i}-\beta^{x}_{i-1},\rho_{i}=1+\beta^{x}_{i}$ given the schedule $\beta^{x}_{i}=\sigma_{1}^{-2i/n}-1$ for a positive $\sigma_{1}$ and $n$ steps.